OpenStudy (anonymous):

Pleaseee someone help me with this. I just can't understand this lesson!! Provide the ordered pairs for two similar triangles on the coordinate plane. Explain, using complete sentences, how you know these triangles are similar.

OpenStudy (tkhunny):

(0,0)(1,0)(0,1) (0,0)(2,0)(0,2) Reason 1) Everything is proportional. Reason 2) AAA I'll let you worry about the Sentence Structure.

OpenStudy (anonymous):

Thank you so much ! can you help me with one other one please ??

OpenStudy (anonymous):

Provide the ordered pairs for two triangles that have no relationship on the coordinate plane. Explain, using complete sentences, how you know these triangles have no relationship

OpenStudy (tkhunny):

Just move one point on one of them. (0,2) ==> (0,6) 1) Is everything still proportional? 2) Do we have AAA?

OpenStudy (anonymous):

(0,0)(1,0)(0,1) (0,0)(2,0)(0,6) so those would be the ordered pairs? Because all 3 aren't similar?

OpenStudy (tkhunny):

You have three sides each, for two traingles. What are the ratios of the corresponding sides? (0,0)(1,0)==> (0,0)(2,0) Ratio is 1:2 (0,0)(0,2)==> (0,0)(0,6) Ratio is 1:3 Want to do the third side? You'll need the Pytharogean Theorem.

OpenStudy (anonymous):

wouldn't it be (0.0)(0,1) = (0,0)0,6) then the third would be a^2 + b^2 = c^2? but what would be a and b?

OpenStudy (tkhunny):

?? We did the sides ON the x- and y-axes. Now do the hypotenuses. (1,0)(0,1) is \[\sqrt{2}\] (2,0)(0,6) is \[\sqrt{40} = 4\cdot\sqrt{10} = 4\cdot\sqrt{5}\cdot\sqrt{2}\] Ratio is \[1:4\cdot\sqrt{5}\] Which isneither 1:2 nor 1:3 -- Definitely NOT "Similar".